# Quantum Steenrod operations of symplectic resolutions

### v2

@@ 3.2.1. Compactly intersecting incidence constraints
We will show that the quantum Steenrod operations can be defined on the submodule of $H^{*}_{\mathbf{T}}(X)$ spanned by the image of Poincaré duals of conormal cycles $[Y_{w}]$, by showing the compactness and transversality of the moduli spaces with incidence constraints on the conormal cycles which define the structure constants. The following is the key lemma.
@@
Recall that $\mathbb{Y}$ is defined as the union of the conormal cycles $Y_{w}=\bigsqcup_{v\leq w}N^{*}\mathring{X}_{v}$. The Schubert cells $\mathring{X}_{v}$ are defined as stable manifolds for the Morse function induced by the (real) moment map $G/B\to\mathbb{R}$ induced by the choice of the dominant cocharacter $\beta:S^{1}\to T$. If one considers the corresponding dominant *$\mathbb{C}^{*}\leq T_{\mathbb{C}}$-action acting on $T^{*}(G/B)$, the conormal bundle $N^{*}\mathring{X}_{w}$ can therefore be identified with all points in $T^{*}(G/B)$ that limit to the fixed point $w$ under the dominant $\mathbb{C}^{*}$-action.
Consider the complex moment map $\mu:T^{*}(G/B)\to\mathfrak{g}^{*}$ which is identified with the Springer resolution, 3.2, with the embedding of the image $\mathcal{N}\subseteq\mathfrak{g}^{*}$. Note that $\mathfrak{g}^{*}$ is a $T$-representation and the moment map $\mu$ is $T$-equivariant. Decompose $\mathfrak{g}^{*}=V_{+}\oplus V_{0}\oplus V_{-}$ for $\beta$-positive, $\beta$-null and $\beta$-negative $T$-weight spaces $V_{+}$, $V_{0}$ and $V_{-}$. By equivariance, the union $\mathfrak{f}_{+}$ of conormals at infinity is projected under $\mu$ to the ideal boundary of $V_{+}$. Similarly, the opposite conormals $\mathfrak{f}_{-}$ at infinity is projected to the ideal boundary of $V_{-}$. Since the ideal boundaries of $V_{+}$ and $V_{-}$ are disjoint, the intersection of the images of $Y_{w}$ and $Y^{v}$ under $\mu$ must intersect in a compact subset. ∎
@@ 3.2.2. Maximum principle
For a cohomology class $b\in H^{*}_{\mathbf{T}}(X;\mathbb{F}_{p})$ Poincaré dual to a ${\mathbf{T}}$-invariant locally finite homology cycle $[Y]$ and take $b_{0}=\mathrm{PD}[Y_{w}]\in H^{*}_{\mathbf{T}}(X;\mathbb{F}_{p})$ and $b_{\infty}=\mathrm{PD}[Y^{v}]\in H^{*}_{\mathbf{T}}(X;\mathbb{F}_{p})$ for $v,w\in W$. Take the product $\mathcal{Y}=Y_{w}\times Y^{\times p}\times Y^{v}\subseteq X\times X^{p}\times X$ and the corresponding incidence cycle $\mathcal{Y}^{eq}$.
@@
Note that the moment map $\mu:T^{*}(G/B)\to\mathcal{N}$ is holomorphic, and by maximum principle applied to the affine variety $\mathcal{N}$, the holomorphic curves in $T^{*}(G/B)$ can only occur in the fiber of $\mu$. Hence the curves with incidence constraints on $Y_{w}$ and $Y^{v}$ lie in the preimage of $\mu(Y_{w})\cap\mu(Y^{v})\subseteq\mathcal{N}$. By 3.11 this subset is contained in a compact subset $K\subseteq\mathcal{N}$, and the result follows for the subset $\mu^{-1}(K)$ from the properness of $\mu$. The bubbles must also lie in the fibers of $\mu$, and their image is connected to the image of the principal component, hence the bubbles are also contained in the same compact subset $\mu^{-1}(K)$. ∎
Hence without any introduction of perturbation of the almost complex structure or inhomogeneous terms, the moduli space is indeed compact. Using compactness of the genuine holomorphic curves, the regularity of the moduli space of solutions to the perturbed Cauchy–Riemann equation can be shown as follows.
@@
For a generic choice of $(J^{eq},\nu^{eq})$, the moduli space with incidence constraints $\mathcal{M}_{A}^{eq;i;2j}\pitchfork\mathcal{Y}^{eq;i;2j}$ (cf. (2.36)) is regular. Moreover, the image of solutions are contained in a compact subset of $X$ (and hence Gromov compactness applies).
@@
From 3.13, there exists a compact subset $K_{0}\subseteq X$ such that all holomorphic curves satisfying the incidence constraints lie in $K_{0}$. Fix a bounded open subset $U\supseteq K_{0}$ and consider only the perturbation data $(J,\nu)$ supported in $U$. For applying standard transversality and compactness results, it suffices to show that there is a compact subset $K_{1}\supseteq U$ only depending on $U$ such that any solution for the perturbation data $(J,\nu)$ such that $\mathrm{supp}(J,\nu)\subseteq U$ must be contained in $K_{1}$.
Let $u:C\to X$ be a solution of ¯$\overline{\partial}_{J}u=\nu$ satisfying the incidence constraints, and consider the projection $\mu\circ u:C\to\mathcal{N}$. By the support condition on $(J,\nu)$, the curve $\mu\circ u$ outside $\mu(U)$ is an unperturbed holomorphic curve. By maximum principle this implies that $\mu\circ u$ is a constant curve outside $\mu(U)$, and $u$ is contained in the fiber of $\mu$ outside $\mu(U)$; this contradicts 3.13. Hence a priori there is a compact subset $\mathcal{N}\supseteq K^{\prime}\supseteq\mu(U)$ such that all solutions $u$ of the perturbed equation satisfy $\mu(\mathrm{im}(u))\subseteq K^{\prime}$. Setting $K_{1}=\mu^{-1}(K^{\prime})$, which is compact by properness of $\mu$, yields the desired compact subset of $X$.
@@
By 3.7 and the remark following the definition, to define $\Sigma_{b}^{\mathbf{T}}$ on the localized equivariant cohomology it suffices to define it on the stable envelope classes, which are in turn linear combinations of the classes represented by the conormal cycles. When the incidence constraints at $0,\infty\in C$ are given by conormal cycles, 3.14 shows that the corresponding moduli spaces required to define the quantum Steenrod operations on the submodule generated by the conormal cycles are indeed regular, hence the operation is well-defined. ∎
@@ 3.3. Regularity of moduli spaces for shift operators
We now show that the shift operators can also be defined on $X=T^{*}(G/B)$, see 2.23. For necessary compactness properties of the moduli spaces to hold, we must assume a condition for the cocharacter (see 3.16 below). To lift this assumption on cocharacter for the definition of shift operators, one must pass to localized equivariant cohomology and use the definition of shift operators via localization from the algebro-geometric construction.
@@
Denote the weights of ${\mathbf{T}}=S^{1}\times T=S^{1}\times(S^{1})^{r}$ by $h=\lambda_{0},\lambda_{1},\dots,\lambda_{r}$. Decompose $\mathfrak{g}^{*}$ into its ${\mathbf{T}}$-weight spaces,
(3.11) $\mathfrak{g}^{*}=\bigoplus_{i=1}^{\dim_{\mathbb{C}}\mathfrak{g}^{*}}\mathbb{C}[f_{i}],\quad f_{i}:=f_{i}(h,\lambda)=\sum_{j=0}^{r}n_{ij}\lambda_{j},\ n_{ij}\in\mathbb{Z}.$
Then 3.16 translates to the condition that $\langle\beta,f_{i}\rangle=\sum_{j=0}^{r}n_{ij}\beta_{j}\geq 0$ for all $i=1,\dots,\dim\mathfrak{g}^{*}$. Consider the fibration $\mathcal{O}(-1)\otimes_{\beta}\mathfrak{g}^{*}\to\mathbb{P}^{1}$ with fibers isomorphic to $\mathfrak{g}^{*}$, where $\mathfrak{g}^{*}$ is glued in to the principal $S^{1}$-bundle over $\mathbb{P}^{1}$ using its ${\mathbf{T}}$-action and the cocharacter $\beta$. Then the complex moment map $\mu:X\to\mathfrak{g}^{*}$ applied fiberwise over $\mathbb{P}^{1}$ yields a holomorphic map
@@
It suffices to show that there is a compact subset of $\mathcal{V}$ which contains all holomorphic sections of $\mathcal{V}\to\mathbb{P}^{1}$ with incidence constraints given by $\mu(Y_{w})$ and $\mu(Y^{v})$. The only nonzero holomorphic sections can arise for the subbundle spanned by summands such that $\langle\beta,f_{i}\rangle=\sum_{j=0}^{r}n_{ij}\beta_{j}=0$, which forms a trivial vector bundle over $\mathbb{P}^{1}$. Considering a nonzero section of this subbundle as a graph of a constant map $p:C\to\mathfrak{g}^{*}$, note that for the section to satisfy the incidence constraints the image of $p$ must lie in the intersection $\mu(Y_{w})\cap\mu(Y^{v})$, which is compact by 3.11. ∎
@@
Fix a cocharacter $\beta:S^{1}\to{\mathbf{T}}$ satisfying 3.16. For a generic choice of $(\widetilde{J}^{eq},\nu^{eq}_{E})$, the moduli space with incidence constraints $\mathcal{M}_{A}^{eq;i;2j}(E_{\beta})\pitchfork\mathcal{Y}^{eq;i;2j}(E_{\beta})$ is regular. Moreover, the image of solutions are contained in a compact subset of $E_{\beta}$.
@@
The proof is essentially the same as in 3.14. Via 3.17, one can choose a compact subset of $E_{\beta}$ containing all holomorphic curves satisfying the incidence constraints, and consider perturbation data $(\widetilde{J},\nu_{E})$ supported in that bounded region. One applies the fiberwise moment map $\widetilde{\mu}:E_{\beta}\to\mathcal{V}=\mathcal{O}(-1)\otimes_{\beta}\mathfrak{g}^{*}$ (see the notation from 3.17), and apply maximum principle in $\mathcal{V}$ to obtain an a priori bound for the image of perturbed solutions under $\widetilde{\mu}$. By properness of $\widetilde{\mu}$, the preimage gives the desired compact subset of $E_{\beta}$ containing all perturbed solutions. Then standard transversality techniques and Gromov compactness apply. ∎
Corollary 3.19 (cf. 3.15).
@@
For later purposes, it is important to consider the inverses of the shift operators $\mathbb{S}_{\beta}$, which may be defined as the reverse shift operators $\mathbb{S}_{-\beta}$; however, the cocharacters $-\beta$ would not satisfy nonnegativity condition, 3.16. Indeed, the structure constants of the inverse would involve poles in the equivariant parameters due to the non-compactness of the relevant moduli spaces. To bypass the need to discuss invertibility separately we will impose the following assumption in our discussion. The following assumption is essentially a comparison of algebro-geometric and symplecto-geometric definitions of Gromov–Witten invariants, and for semipositive targets such as $X$ it is expected to be true.

### v3

@@ 3.2.1. Compactly intersecting incidence constraints
Following the remark after 3.10, we will show that the quantum Steenrod operations can be defined on the submodule of $H^{*}_{\mathbf{T}}(X)$ spanned by the image of Poincaré duals of conormal cycles $[Y_{w}]$. We must show the compactness and transversality of the moduli spaces with incidence constraints on the conormal cycles which define the structure constants.
The difference from the previous general setup is that on the submodule of $H^{*}_{\mathbf{T}}(X)$ spanned by the conormal cycles, we have a control of non-compactness of the incidence cycles at $0\in C$, as the conormal cycles approach infinity at prescribed directions (the conormal directions to the stable manifolds $\mathring{X}_{w}$). Consequently, we are not restricted to considering only the compact cycles as incidence constraints at $\infty\in C$, but we can consider locally finite cycles as long as they properly intersect the conormal cycles. The following is the key lemma for establishing this.
@@
Recall that $\mathbb{Y}$ is defined as the union of the conormal cycles $Y_{w}=\bigsqcup_{v\leq w}N^{*}\mathring{X}_{v}$. The Schubert cells $\mathring{X}_{v}$ are defined as stable manifolds for the Morse function induced by the (real) moment map $G/B\to\mathbb{R}$ induced by the choice of the dominant cocharacter $\beta:S^{1}\to T$. If one considers the corresponding dominant $\mathbb{C}^{*}\leq T_{\mathbb{C}}$-action acting on $T^{*}(G/B)$, the conormal bundle $N^{*}\mathring{X}_{w}$ can therefore be identified with all points in $T^{*}(G/B)$ that converges to the fixed point $w$ under the dominant $\mathbb{C}^{*}$-action (as $z\in\mathbb{C}^{*}$ approaches $z\to\infty$).
Consider the complex moment map $\mu:T^{*}(G/B)\to\mathfrak{g}^{*}$ which is identified with the Springer resolution, 3.2, with the embedding of the image $\mathcal{N}\subseteq\mathfrak{g}^{*}$. Note that $\mathfrak{g}^{*}$ is a $T$-representation and the moment map $\mu$ is $T$-equivariant. Decompose $\mathfrak{g}^{*}=V_{+}\oplus V_{0}\oplus V_{-}$ for $\beta$-positive, $\beta$-null and $\beta$-negative $T$-weight spaces $V_{+}$, $V_{0}$ and $V_{-}$. By equivariance, the union of ideal boundary of $\mathbb{Y}$ is projected under $\mu$ to the ideal boundary of $V_{+}$. Similarly, the opposite conormals $\mathbb{Y}^{-}$ at infinity is projected to the ideal boundary of $V_{-}$. Since the ideal boundaries of $V_{+}$ and $V_{-}$ are disjoint, the intersection of the images of $Y_{w}$ and $Y^{v}$ under $\mu$ must intersect in a compact subset. ∎
@@ 3.2.2. Maximum principle
Fix a cohomology class $b\in H^{*}_{\mathbf{T}}(X;\mathbb{F}_{p})$ Poincaré dual to a $\mathbf{T}$-invariant locally finite homology cycle $[Y]$ and take $b_{0}=\mathrm{PD}[Y_{w}]\in H^{*}_{\mathbf{T}}(X;\mathbb{F}_{p})$ and $b_{\infty}=\mathrm{PD}[Y^{v}]\in H^{*}_{\mathbf{T}}(X;\mathbb{F}_{p})$ for $v,w\in W$. Take the product $\mathcal{Y}=Y_{w}\times Y^{\times p}\times Y^{v}\subseteq X\times X^{p}\times X$ and the corresponding incidence cycle $\mathcal{Y}^{eq}$.
@@
Note that maximum principle applies to the affine variety $\mathcal{N}\subseteq\mathfrak{g}^{*}$, and the holomorphic curves in $\mathcal{N}$ must be constant. The moment map $\mu:T^{*}(G/B)\to\mathcal{N}$ is holomorphic, so the holomorphic curves in $T^{*}(G/B)$ can only occur in the fiber of $\mu$. Hence the curves with incidence constraints on $Y_{w}$ and $Y^{v}$ lie in the preimage of $\mu(Y_{w})\cap\mu(Y^{v})\subseteq\mathcal{N}$. By 3.13 this subset is contained in a compact subset $K\subseteq\mathcal{N}$, and the result follows for the subset $\mu^{-1}(K)$ from the properness of $\mu$. The bubbles must also lie in the fibers of $\mu$, and their image is connected to the image of the principal component, hence the bubbles are also contained in the same compact subset $\mu^{-1}(K)$. ∎
Hence without any introduction of perturbation of the almost complex structure or inhomogeneous terms, the moduli space is indeed compact. Using compactness of the genuine holomorphic curves, the regularity of the moduli space of solutions to the perturbed Cauchy–Riemann equation can be shown as follows. As usual in Floer theory, we prove regularity of the moduli spaces for generic choices of perturbation data; in our case, the equivariant perturbation data are chosen from
(3.8) $J^{eq}$ =$=\{J_{w^{\prime}}\}_{w^{\prime}\in ET}\in C^{\infty}(ET\times X;\mathrm{End}(TX)),$
(3.9) $\nu^{eq}$ =$=\{\nu_{w,w^{\prime}}\}\in C^{\infty}(E\mathbb{Z}/p\times ET\times C\times X;\mathrm{Hom^{0,1}}(TC;TX)).$
Fix the integrable perturbation data $(J^{eq}_{0},\nu^{eq}_{0})=(J,0)$ where $J$ is the (constant over $ET$) integrable complex structure, and consider its neighborhood $\mathcal{S}\ni(J^{eq}_{0},\nu^{eq}_{0})$ in the above spaces (in the $C^{\infty}$ topology).
@@
There exists a comeager subset $\mathcal{S}^{\mathrm{reg}}\subseteq\mathcal{S}$ of equivariant perturbation data $(J^{eq},\nu^{eq})$ near $(J^{eq}_{0},\nu^{eq}_{0})$, such that for any $(J^{eq},\nu^{eq})\in\mathcal{S}^{\mathrm{reg}}$, the moduli space with incidence constraints $\mathcal{M}_{A}^{eq;i;2j}\pitchfork\mathcal{Y}^{eq;i;2j}$ (cf. (2.38)) is regular. Moreover, the image of solutions are contained in a compact subset of $X=T^{*}(G/B)$ (and hence Gromov compactness applies).
@@
From 3.15, there exists a compact subset $K_{0}\subseteq X$ such that all holomorphic curves satisfying the incidence constraints lie in $K_{0}$. Fix a bounded open subset $U\supseteq K_{0}$ and consider only the perturbation data $(J,\nu)$ supported in $U$. For applying standard transversality and compactness results, it suffices to show that there is a compact subset $K_{1}\supseteq U$ only depending on $U$ such that any solution for the perturbation data $(J,\nu)$ such that $\mathrm{supp}(J,\nu)\subseteq U$ must be contained in $K_{1}$.
Let $u:C\to X$ be a solution of ¯$\overline{\partial}_{J}u=\nu$ satisfying the incidence constraints, and consider the projection $\mu\circ u:C\to\mathcal{N}$. We show that there exists a compact subset in $\mathcal{N}$ which contains the image of $\mu\circ u$ for all solutions $u$. Suppose otherwise, and take a curve $u$ such that $\mu\circ u$ has a nonzero intersection with the bounded subset $\mu(U)\subseteq\mathcal{N}$. By the support condition on $(J,\nu)$, the curve $\mu\circ u$ outside $\mu(U)$ is an unperturbed holomorphic curve. By maximum principle this implies that $\mu\circ u$ is a constant curve outside $\mu(U)$, and $u$ is contained in the fiber of $\mu$ outside $\mu(U)$. This contradicts 3.15, since by assumption $u$ must be contained in the fiber of $\mu$ over $\mu(K_{0})\subseteq\mu(U)$. Hence a priori there is a compact subset $\mathcal{N}\supseteq K^{\prime}\supseteq\mu(U)$ such that all solutions $u$ of the perturbed equation satisfy $\mu(\mathrm{im}(u))\subseteq K^{\prime}$. Setting $K_{1}=\mu^{-1}(K^{\prime})$, which is compact by properness of $\mu$, yields the desired compact subset of $X$.
@@
Remark 3.17.
Generic transversality statements do not directly apply to spaces of $C^{\infty}$-sections; rather, one (i) first passes to a suitable Banach completion (such as Floer’s $C^{\infty}_{\varepsilon}$-spaces) to which one applies the generic transversality results, and (ii) uses e.g. the Taubes trick (see [Wen, Section 4.4.2]) to pass to genericity in the $C^{\infty}$-topology. ∎
@@
By 3.8, 3.10 and the remarks following them, to define $\Sigma_{b}^{\mathbf{T}}$ on the localized equivariant cohomology it suffices to define it on the stable envelope classes, which are in turn linear combinations of the classes represented by the conormal cycles. When the incidence constraints at $0,\infty\in C$ are given by conormal cycles, 3.16 shows that the corresponding moduli spaces required to define the quantum Steenrod operations on the submodule generated by the conormal cycles are indeed regular, hence the operation is well-defined. ∎
@@ 3.3. Regularity of moduli spaces for shift operators
We now show that the shift operators can also be defined on $X=T^{*}(G/B)$, see 2.23. For necessary compactness properties of the moduli spaces to hold, we must assume a condition for the cocharacter (see 3.19 below). To lift this assumption on cocharacter for the definition of shift operators, one must pass to localized equivariant cohomology and use the definition of shift operators via localization from the algebro-geometric construction.
@@
Denote the weights of $\mathbf{T}=S^{1}\times T=S^{1}\times(S^{1})^{r}$ by $h,\lambda_{1},\dots,\lambda_{r}$. Decompose $\mathfrak{g}^{*}$ into its $\mathbf{T}$-weight spaces,
(3.12) $\mathfrak{g}^{*}=\bigoplus_{i=1}^{\dim_{\mathbb{C}}\mathfrak{g}^{*}}\mathbb{C}[f_{i}],\quad f_{i}:=n_{i0}h+\sum_{j=1}^{r}n_{ij}\lambda_{j}\in(\mathrm{Lie}\mathbf{T})^{*},\ n_{ij}\in\mathbb{Z}.$
Denoting the components of $\beta$ by $\langle\beta,\lambda_{j}\rangle=\beta_{j}$, 3.19 translates to the condition that $\langle\beta,f_{i}\rangle=\sum_{j=0}^{r}n_{ij}\beta_{j}\geq 0$ for all $i=1,\dots,\dim\mathfrak{g}^{*}$. Consider the associated bundle $\mathcal{O}_{\mathbb{P}^{1}}(-1)\otimes_{\beta}\mathfrak{g}^{*}\to\mathbb{P}^{1}$ with fibers isomorphic to $\mathfrak{g}^{*}$, defined as
(3.13) $\mathcal{O}_{\mathbb{P}^{1}}(-1)\otimes_{\beta}\mathfrak{g}^{*}:=\{([z],y)\in\mathcal{O}(-1)\times\mathfrak{g}\}/([z],y)\sim([z]\cdot\theta,\beta(\theta^{-1})\cdot y)\mbox{ for }\theta\in S^{1}$
using the cocharacter $\beta:S^{1}\to\mathbf{T}$. Then the complex moment map $\mu:X\to\mathfrak{g}^{*}$ applied fiberwise over $\mathbb{P}^{1}$ yields a holomorphic map
@@
It suffices to show that there is a compact subset of $\mathcal{V}$ which contains all holomorphic sections of $\mathcal{V}\to\mathbb{P}^{1}$ with incidence constraints given by $\mu(Y_{w})$ and $\mu(Y^{v})$. The only nonzero holomorphic sections can arise for the subbundle spanned by summands such that $\langle\beta,f_{i}\rangle=\sum_{j=0}^{r}n_{ij}\beta_{j}=0$, which forms a trivial vector bundle over $\mathbb{P}^{1}$. Considering a nonzero section of this subbundle as a graph of a constant map $p:C\to\mathfrak{g}^{*}$, note that for the section to satisfy the incidence constraints the image of $p$ must lie in the intersection $\mu(Y_{w})\cap\mu(Y^{v})$, which is compact by 3.13. The result follows from the properness of $\mu$ (it is the Springer resolution, see 3.2). ∎
Below, as in the proof of 3.16, we consider equivariant perturbation data $(\widetilde{J}^{eq},\nu_{E}^{eq})$ in the neighborhood $\mathcal{S}$ (in the $C^{\infty}$-topology) of the reference perturbation data $(\widetilde{J}^{eq}_{0},\nu_{E,0}^{eq})=(\widetilde{J},0)$. A generic choice is interpreted as a choice in a comeager subset of such neighborhood.
@@
Fix a cocharacter $\beta:S^{1}\to\mathbf{T}$ satisfying 3.19. For a generic choice of $(\widetilde{J}^{eq},\nu^{eq}_{E})$, the moduli space with incidence constraints $\mathcal{M}_{A}^{eq;i;2j}(E_{\beta})\pitchfork\mathcal{Y}^{eq;i;2j}(E_{\beta})$ is regular. Moreover, the image of solutions are contained in a compact subset of $E_{\beta}$.
@@
The proof is essentially the same as in 3.16. Via 3.20, one can choose a compact subset of $E_{\beta}$ containing all holomorphic curves satisfying the incidence constraints, and consider perturbation data $(\widetilde{J},\nu_{E})$ supported in that bounded region. One applies the fiberwise moment map $\widetilde{\mu}:E_{\beta}\to\mathcal{V}=\mathcal{O}(-1)\otimes_{\beta}\mathfrak{g}^{*}$ (see the notation from 3.20), and apply maximum principle in $\mathcal{V}$ to obtain an a priori bound for the image of perturbed solutions under $\widetilde{\mu}$. By properness of $\widetilde{\mu}$, the preimage gives the desired compact subset of $E_{\beta}$ containing all perturbed solutions. Then standard transversality techniques and Gromov compactness apply. ∎
Corollary 3.22 (cf. 3.18).
@@
For later purposes, it is important to consider the inverses of the shift operators $\mathbb{S}_{\beta}$, which may be defined as the reverse shift operators $\mathbb{S}_{-\beta}$; however, the cocharacters $-\beta$ would not satisfy nonnegativity condition, 3.19. Indeed, the structure constants of the inverse would involve poles in the equivariant parameters due to the non-compactness of the relevant moduli spaces. To bypass the need to discuss invertibility separately we will impose the following assumption in our discussion. The following assumption is essentially a comparison of algebro-geometric and symplecto-geometric definitions of Gromov–Witten invariants. For semipositive targets such as $X$, it is expected to be true.